# OmniVLA: Physically-Grounded Multimodal VLA with Unified Multi-Sensor Perception for Robotic Manipulation

### v2

@@ Abstract
Vision-language-action (VLA) models have shown strong generalization for robotic action prediction through large-scale vision-language pretraining. However, most existing models rely solely on RGB cameras, limiting their perception and, consequently, manipulation capabilities.
We present OmniVLA, an omni-modality VLA model that integrates novel sensing modalities for physically-grounded spatial intelligence beyond RGB perception. The core of our approach is the sensor-masked image, a unified representation that overlays spatially grounded and physically meaningful masks onto the RGB images, derived from sensors including an infrared camera, a mmWave radar, and a microphone array. This image-native unification keeps sensor input close to RGB statistics to facilitate training, provides a uniform interface across sensor hardware, and enables data-efficient learning with lightweight per-sensor projectors. Built on this, we present a multisensory vision-language-action model architecture and train the model based on an RGB-pretrained VLA backbone. We evaluate OmniVLA on challenging real-world tasks where sensor-modality perception guides the robotic manipulation. OmniVLA achieves an average task success rate of 84%, significantly outperforms both RGB-only and raw-sensor-input baseline models by 59% and 28% respectively, meanwhile showing higher learning efficiency and stronger generalization capability.
@@ I Introduction
Vision–language–action (VLA) models [1, 2] recently emerged as a powerful paradigm towards generalist policies for embodied AI. They leverage vision-language pretraining to interpret user prompts and camera observations for physical system control, including robots, autonomous vehicles, and even factories and warehouse facilities. Although showing great generalization capability and instruction following, most VLA models are limited to RGB camera image input, which constrains their perception capabilities, preventing them from handling tasks that require non-RGB cues. This undermines the potential of robots to utilize additional sensory hardware and perform challenging tasks that require perception capability similar to or even beyond humans. For example, Infrared (IR) cameras reveal temperature contrast for tasks such as search and rescue; millimeter wave (mmWave) radars penetrate occlusions, such as cardboard or clothing, with radio-frequency (RF) signals to localize hidden objects; acoustic microphone arrays enable human-like environmental awareness to ambient sound and react accordingly. Adding per-sensor encoders or simply feeding raw sensor data to VLA models, however, increases system complexity, suffers from scarce training data available, and weakens compatibility with vision-language pretraining.
Figure 1: Instead of relying solely on RGB cameras, OmniVLA equips embodied AI with multi-sensor perception beyond-RGB. We use beamforming to construct heatmap-like sensor images for acoustic and mmWave to highlight the sound source and the hidden item, respectively.
In this paper, we aim to enable VLA models to perceive multi-sensor data input, utilizing various sensor modalities including infrared, mmWave, and acoustic. This enables robots to combine the strong generalization of foundation models and physical information from various sensors seamlessly to enable physically-grounded spatial intelligence.
There are several challenges in integrating diverse sensors with a VLA model. First, VLA models need to effectively interpret heterogeneous sensor information and use that to guide the action output; naively feeding or stacking raw sensor streams provides poor performance and data efficiency, as shown in Section IV, because existing VLA backbones are trained on RGB image data. Second, sensors differ in format, field of view, and resolution, calling for a scalable, uniform representation rather than training sensor fusion models that depend on specific hardware. Lastly, sensor modality data are much more scarce compared to web-scale image-text data pairs, so we require a data-efficient approach for training. To solve these challenges, we take inspiration from how the human brain interprets sensor information: as we are used to RGB images, we naturally anchor other sensor cues to the RGB view—for example, interpreting infrared camera images by associating temperatures with objects visible in the image.
@@
Built on the sensor-masked images, we propose a tailored VLA model architecture (Figure 2). We first convert all raw sensor measurements to image-like 2D spatial representations. Specifically, we perform beamforming for mmWave radar and acoustic array data to acquire heatmap-like raw sensor images. Then, to generate masks for interested objects in the scene, we use a cloud-based Vision-Language Model (VLM) to interpret the task request and generate a prompt for grounded SAM to provide semantic-based segmentation masks [3]. Note that the VLM is only invoked once per task, so its delay does not affect the real-time robot actions. We further overlay sensor information on the RGB images in masked regions to acquire the sensor-masked images as the input to the frozen vision encoder. After the vision encoder, we add lightweight projection layers for each sensor to generate better-aligned tokens for sensor images. Finally, the input tokens are processed by an LLM, which provides tokens to the action expert for the robot action output.
We build a multi-sensor robot arm prototype to collect RGB camera data together with various sensor data, paired with expert demonstrations, and train OmniVLA with the collected datasets. We evaluate OmniVLA extensively with several manipulation tasks that require sensor-modality guidance, including thermal-based pick-and-place, mmWave-based see-through boxes and opening the non-empty one, and acoustic-based uncovering a ringing phone beneath clothes. OmniVLA achieves an average task success rate of 84%, significantly outperforms 25% success rate of RGB-only baseline and 56% success rate of raw-sensor-input baseline. This shows the benefit from our unique sensing capabilities and the performance gain from our sensor-masked image representation. We also highlight the data efficiency of our approach by achieving a similar success rate with only 50% training data compared to the raw-sensor-input baseline. Moreover, we show that our approach provides strong generalization capability across three unseen tasks, outperforming the base VLA model and raw-sensor-data-based model by 59% and 28% respectively on average success rate.
@@
    To our knowledge, OmniVLA is the first VLA model that unifies multiple sensing modalities, including infrared, mmWave, and acoustic, to enable robotic manipulation tasks that exceed RGB-based perception capability.
@@
    We introduce sensor-masked images, a spatially grounded and semantically aligned representation that allows the reuse of a pre-trained vision encoder, provides a uniform representation across sensor hardware, and improves task learning efficiency.
@@
    We present a lightweight OmniVLA model architecture and evaluate system model performance with extensive experimental evaluation.
@@ II-A Vision-Language-Action Models
Vision-language-action (VLA) models have been a popular research paradigm for robotic manipulation, using the language prompt and video feed as input and generating robot actions in an end-to-end manner. Conventional robotic manipulation policies using reinforcement learning and simulators provide great performance on specific tasks [4, 5, 6, 7, 8, 9, 10], while VLA models have shown great improvement in few-shot task generalization and instruction following by leveraging web-scale pretraining. The majority of VLA models only take video from RGB cameras as visual input [11, 12, 13, 14, 1, 2]. However, these works are inherently limited by the RGB camera input and are unable to finish tasks that require perception capabilities beyond RGB.
To address this issue, researchers propose novel Vision-Language-Action models with additional sensor input. Depth information is widely employed to enhance the capability of Vision-Language-Action models with better spatial-temporal understanding [15, 16, 17, 18, 19]. Other works incorporate tactile perception in Vision-Language-Action models for better task planning, grounding, and reasoning capabilities [20, 21]. Moreover, Vlas [22] integrates speech information into Vision-Language-Action models for convenient and personalized human-robot interaction. MultiPLY [23] introduces a embodied LLM for planning multisensory interactions with the environment, but it is based on simulator only and cannot perform detailed action control.
Prior work has not explored the integration of novel sensing modalities like thermal, mmWave, and acoustic, which provide unique sensing information for manipulation tasks. Also, existing approaches often require training sensor encoders for each modality, which need a large amount of data. In contrast, we propose a data-efficient approach towards beyond-RGB perception. Moreover, previous works propose complex and specific model architectures for extra sensor and image fusion, which are not generalizable to diverse sensors. In contrast, we propose Sensor-Masked Image, a simple and unified sensor fusion strategy for Vision-Language-Action models.
@@ II-B Multi-sensor Fusion based Perception
Multi-sensor fusion based perception has received significant interest recently, especially for 3D detection and other downstream tasks for autonomous driving. The benefits for multi-sensor fusion are comprehensive environmental understanding for complementary sensors [24]. One typical type of multi-sensor fusion is mmWave radar and camera sensor fusion, widely used in autonomous driving because of its reliability in rainy and foggy environments [25, 26, 27, 28, 29, 30, 31]. Another thread of recent work fuses RF signals for robotic perception to achieve non-line-of-sight object finding [32, 33, 34], but it requires extra RFID tags attached to the objects and is specifically designed for RF signals.
@@ III System Design
Figure 2: System Overview. OmniVLA processes diverse sensor data into image-like 2D spatial representations, overlaying sensor information on top of RGB images to acquire spatially grounded and semantically aligned masked-sensor images. We train a VLA model with individual MLP sensor projectors to achieve challenging tasks requiring beyond-RGB perception.
@@ III-A System Overview
OmniVLA contains two parts: sensor-masked image generation and multi-sensor vision-language-action model architecture. The first part first extracts raw sensor data, then preprocesses it into 2D sensor images/heatmaps. After that, we use a semantic-based segmentation model to generate the masks from the RGB images following a VLM-generated prompt. Then we overlay the sensor images on the masked regions of RGB images to output sensor-masked images, which are the input for our multi-sensor vision-language-action model. The second part is a multi-sensor vision-language-action model backbone that is designed for sensor-masked images to capture each sensor’s input and avoid requiring a large amount of sensor data for training. We utilize the existing frozen vision encoders to encode sensor-masked images. For each sensor modality, we use individual multi-layer perceptron layers to align sensor image tokens with language and RGB image tokens. The tokens are concatenated together with language tokens as input for the large language model in the architecture, and then we generate the final action predictions using the action expert module.
Figure 3: Sensor Data Processing Illustration. We propose a general sensor data processing pipeline applicable to various sensors, including (a) thermal camera, (b) mmWave radar, and (b) acoustic microphone array, by overlaying sensor information on top of RGB images as VLA model input. We update prompt input to SAM2 model when the task begins and then asynchronously check for updates, so that VLM output delay does not affect the real-time processing of sensor-masked images.
@@ III-B Sensor Data Representation
The first part of our work is sensor-masked image generation, which includes sensor data preprocessing, segmentation mask processing, sensor-masked image blending.
Preprocessing. We convert all raw measurements into a camera-like 2D spatial representation. Thermal camera already outputs raster images—infrared intensity—defined over image coordinates $(u,v)$. In contrast, the mmWave radar and the microphone array provide complex signal samples per array element, $x_{i,k}=A_{i,k}e^{j\psi_{i,k}}$ for the $k$-th element of array $i$, where $i\in\{\text{mmWave radar},\text{microphone array}\}$. To obtain inputs with consistent spatial mapping (horizontal and vertical viewing axes on a 2D grid), we compute azimuth–elevation heatmaps via conventional delay-and-sum beamforming [37]:
@@
where $\theta$ is the azimuth angle, $\phi$ is the elevation angle, $\lambda_{i}$ is the wavelength received by $i^{th}$ array element (antenna or microphone), $(x_{i,k},y_{i,k})$ is the position for $k^{th}$ array element in array $i$. One example is shown in Fig. 1. Similar to the principle of RGB camera, the azimuth-elevation heatmap reveals the environmental information in a direct way for human understanding and acts as the sensor images in the following steps.
Segmentation and overlay. The next step is generating segmentation masks from the task description and RGB images. We first send the task request text and RGB image to the state-of-the-art vision-language models, GPT-4o [38], to generate the segmentation prompts that describe objects related to the task in the current scene, for example, ‘red block/drink’, ‘black phone’, ‘cardboard boxes’. We update the prompt when the task begins and then asynchronously check for updates, so that VLM output delay does not affect real-time robot manipulation actions. Then we input the segmentation prompts and RGB image into a segmentation model to generate image masks for task-related objects. Specifically, we use Grounded SAM 2 [3], which combines segment anything model [39] and open-set object detection model, grounding DINO [40]. The final output is a 0-1 matrix, where ‘1’ labels the masked regions for objects. The second step can be expressed as the following equations
@@
$\text{mask}=\mathrm{SAM2}(l,\mathrm{I}_{\mathrm{RGB}})$
where $l$ represents the segmentation prompt generated by VLM, $\mathrm{T}_{\text{task}}$ is the input task description, and $\mathrm{I}_{\mathrm{RGB}}$ is the RGB image. The final step is blending sensor images and RGB images to generate sensor-masked images. We perform one-time calibration between the sensors and RGB cameras by rotation and clipping to ensure they are spatially aligned. Note that some inaccuracy is tolerable as we don’t require a precise match. Then, we segment the specific part of sensor images we are interested in by masks in the previous step. Finally, we blend parts of sensor images with RGB images for the final sensor-masked images. The process of the final step is shown in Eq.(3),
@@
where $i\in\{\text{mmWave radar},\text{microphone array},\text{thermal camera}\}$ and $\alpha$ is the hyper-parameter for image blending. The higher $\alpha$ is, the more sensor information remains for comprehensive understanding, but less correlation between the masked part and the unmasked RGB image can be learned from the model. We set $\alpha$ as 1 by default based on our empirical testing. Final sensor-masked images act as input for our multi-sensor Vision-Language-Action model for the second part.
@@ III-C Model architecture and training
The second part of our work is a multi-sensor vision-language-action model for sensor-masked images, which contains the architecture design and training strategy.
In recent Vision-Language-Action model, it first encodes the input RGB images with vision encoder, then projects encoded results with multi-layer perceptron (MLP) layers for image tokens. The image tokens are concatenated with language tokens as input for large language models to further generate actions to manipulate the robot, which we follow to design our architecture.
We present a generalizable and efficient multi-sensory VLA model architecture design as shown in Fig. 2. We enable effective sensor data understanding by feeding sensor-masked images to an existing vision encoder. Then for each sensor modality, we input the encoded results into individual multi-layer perceptron modules for projection to align with the language tokens. The projected tokens for each sensor are concatenated together with language tokens from the task description passing the language encoder. Finally, a large language model takes concatenated tokens as input, then provides tokens or features to a diffusion-based action expert for final robotic action generation. As we focus on integrating sensor information with VLA models for multi-sensory perception, we follow typical model architectures for the action generation, such as SmolVLA [2], pi0 [1]. Our approach is potentially compatible with other VLA models as well due to its image-like sensor data representation.
The overall data flow can be described with the following equations
@@
where $m$ is the number of sensors we use, $\mathrm{E}_{I}$ is the image encoder, and $\mathrm{E}_{L}$ is the language embedding layer, $\text{t}_{i}$ is the embedding of $i$th sensor image. We underline that our architecture does not require all sensors shown here. Instead, we allow flexible sensor setup according to the deployment scenario, budget, etc., for example using a single sensor like an infrared camera. This design is a general and flexible framework to process beyond-RGB sensor-masked images, compatible with existing Vision-Language-Action models. It means we can utilize the vision-language capability in existing model’s pretraining and understand unique features from each sensor easily with a few demonstration data.
During the model training, we freeze the vision and language encoders, typical in various Vision-Language-Action models [2, 1, 41] and Vision-Language models finetuning [42, 43]. Meanwhile, we set all other weights trainable for multi-sensor perception in robotic manipulation. We initialize individual MLP projectors with weights from the RGB projection layer in existing Vision-Language-Action models, which provides a basic understanding of images and allows updating for new sensor image features. We co-finetune MLP modules for each sensor and the rest unfrozen weights of the model using the demonstration dataset.

### v3

@@ Abstract
Vision-language-action (VLA) models have shown strong generalization in robotic manipulation through large-scale vision-language pretraining. However, most existing models rely solely on RGB cameras, limiting their perception and, consequently, manipulation capabilities. We present OmniVLA, an omni-modality VLA model that integrates novel sensing modalities to enable beyond-RGB robotic perception and manipulation. The core of our approach is the sensor-masked image, a unified representation that overlays physically meaningful, spatially grounded masks onto the RGB images. These masks are derived from sensors including an infrared camera, a mmWave radar, and a microphone array. This image-native unification keeps sensor input close to RGB statistics to facilitate training, provides a uniform interface across sensor hardware, and enables data-efficient learning with lightweight per-sensor projectors. Building on this, we design a multimodal vision-language-action model architecture and train OmniVLA by extending an RGB-pretrained VLA backbone. We evaluate OmniVLA on challenging real-world tasks that require sensor-modality perception to guide the manipulation. OmniVLA achieves an average task success rate of 84%, significantly outperforms both RGB-only and raw-sensor-input baseline models by 59% and 28% respectively, meanwhile showing higher learning efficiency and stronger generalization capability.
@@ I Introduction
Vision–language–action (VLA) models [3, 34] recently emerged as a powerful paradigm for generalist robotic policies. They leverage vision-language pretraining to map user prompts and camera observations to robot actions, showing great generalization and instruction following capability. However, most VLA models are limited to RGB camera image input, which constrains their perception capabilities and prevents them from handling tasks that require non-RGB cues. This undermines the potential of robots to utilize additional sensory hardware and perform challenging tasks that require perception capability similar to or even beyond humans. For example, Infrared (IR) cameras reveal temperature contrast for tasks such as search and rescue; millimeter wave (mmWave) radars penetrate occlusions, such as cardboard or clothing, with radio-frequency (RF) signals to localize hidden objects [12, 11]; acoustic microphone arrays enable human-like environmental awareness to ambient sound and react accordingly. Adding per-sensor encoders or directly feeding raw sensor data to VLA models, however, increases system complexity, exacerbates data scarcity, and reduces compatibility with pretrained vision-language models.
Figure 1: Instead of relying solely on RGB cameras, OmniVLA equips robots with multi-sensor perception. We use beamforming heatmaps as acoustic and mmWave sensor images to highlight sound source and hidden item respectively.
In this paper, we aim to equip VLA models with unified multi-sensor perception, utilizing sensor modalities including infrared, acoustic, and mmWave. This enables robots to combine the strong generalization of foundation models and physical information from various sensors seamlessly, to enable physically-grounded spatial intelligence.
There are several challenges in integrating diverse sensors with a VLA model. First, VLA models need to effectively interpret heterogeneous sensor information and use them to guide the action; naively feeding raw sensor streams leads to poor performance and data efficiency as shown in Section IV, because existing VLA backbones are trained primarily on RGB images. Second, sensors differ in format, field of view, and resolution, calling for a scalable, uniform representation rather than training sensor fusion models that depend on specific hardware. Lastly, sensor modality data are much more scarce compared to web-scale image-text data pairs, so we require a data-efficient approach for training. To solve these challenges, we take inspiration from how the human brain interprets sensor information: as we are used to RGB images, we naturally anchor other sensor cues to the RGB view—for example, interpreting infrared camera images by associating temperatures with objects visible in the image.
@@
Building on the sensor-masked images, we propose a tailored VLA model architecture (Figure 2). We first convert all raw sensor measurements to image-like 2D spatial representations, specifically, we perform beamforming for mmWave radar and acoustic array data to acquire heatmap-like raw sensor images. Then, to generate masks for interested objects in the scene, we use a cloud-based Vision-Language Model (VLM) to interpret task request and generate a prompt for grounded SAM [33] to provide semantic-based segmentation masks. We further overlay sensor information on the RGB images at masked regions to acquire the sensor-masked images as the input to a frozen vision encoder. After the vision encoder, we add lightweight projection layers for each sensor to generate better aligned tokens for sensor images. Finally, a LLM backbone processes the input tokens and conditions the diffusion-based action expert for the robot action output.
We build a multi-sensor robot arm prototype to collect RGB camera and sensor data, paired with action demonstrations, and train OmniVLA with the collected datasets. We evaluate OmniVLA extensively with several manipulation tasks that require sensor-modality guidance, including thermal-based pick-and-place, mmWave-based see-through boxes and opening the non-empty one, and acoustic-based uncovering ringing phone beneath clothes. OmniVLA achieves an average task success rate of 84%, significantly outperforms 25% success rate of RGB-only baseline and 56% success rate of raw-sensor-input baseline. This shows the benefit from our unique sensing capabilities, and the performance gain from our sensor-masked image representation. We also highlight the data efficiency of our approach by achieving similar success rate with only 50% training data compared to raw-sensor-input baseline. Moreover, we show that our approach provides strong generalization capability across three unseen tasks, outperforming base VLA model and raw-sensor-data based model by 59% and 28% respectively on average success rate.
@@
    To our knowledge, OmniVLA is the first VLA model that unifies multiple sensing modalities, including infrared, mmWave, acoustic, to enable robotic manipulation tasks beyond RGB-based perception capability.
@@
    We introduce sensor-masked images, a spatially grounded and semantically aligned representation that allows reusing pre-trained vision encoders, provides a uniform representation across sensor hardware, and improves learning efficiency.
@@
    We present a lightweight OmniVLA model architecture and evaluate system performance with extensive experimental evaluation. We open-source the project at https://github.com/GuoHeyu/OmniVLA.
@@ II-A Vision-Language-Action Models
Vision-language-action (VLA) models have been a popular research paradigm for robotic manipulation, using the language prompt and video feed as input and generating robot actions in an end-to-end manner. Conventional robotic manipulation policies using reinforcement learning and simulators provide great performance on specific tasks [9, 27, 40, 21, 8, 20], while VLA models have shown great improvement in few-shot task generalization and instruction following by leveraging web-scale pretraining. The majority of VLA models only take video from RGB cameras as visual input [28, 19, 36, 41, 3, 34]. However, these works are inherently limited by the RGB camera input and unable to finish tasks that require perception capabilities beyond RGB.
To address this issue, researchers propose novel Vision-Language-Action models with additional sensor input. Depth information is widely employed to enhance the capability of VLA models with better spatial-temporal understanding [22, 1, 31, 30, 43]. Other works incorporate tactile perception in VLA models for better task planning, grounding and reasoning capabilities [2, 15]. Moreover, Vlas [42] integrates speech information for convenient and personalized human-robot interaction. MultiPLY [14] introduces an embodied LLM for planning multisensory interactions with the environment, but it is based on simulator only and cannot generate actions for manipulation tasks.
Prior works have not explored the integration of novel sensing modalities like thermal, acoustic, and mmWave, which provide unique sensing information for manipulation tasks. Also, existing approaches often require training sensor encoders for each modality, which needs a large amount of data. In contrast, we propose a data-efficient approach towards beyond-RGB perception. Moreover, previous works propose complex and specific model architecture for extra sensor and image fusion, which is not generalizable to diverse sensors. In contrast, we propose sensor-masked images, a simple and unified sensor fusion strategy for VLA models.
@@ II-B Multi-sensor Fusion based Perception
Multi-sensor fusion based perception has received significant interest recently, especially for 3D detection and other downstream tasks for autonomous driving. The benefits for multi-sensor fusion are comprehensive environmental understanding for complementary sensors [26]. One typical type of multi-sensor fusion is mmWave radar and camera sensor fusion, widely used in autonomous driving because of its reliability in rainy and foggy environments [12, 37, 23, 29, 18, 38]. Another thread of recent work fuses RF signals for robotic perception to achieve non-line-of-sight object finding [6, 5, 4], but it requires extra RFID tags attached on the objects and are designed only for RF signals.
@@ III System Design
Figure 2: System Overview. OmniVLA processes diverse sensor data into image-like 2D spatial representations, and then overlays sensor information onto RGB images to produce spatially grounded and semantically aligned sensor-masked images. We train OmniVLA by extending an RGB-pretrained VLA backbone to perform challenging tasks beyond RGB perception.
@@ III-A System Overview
OmniVLA contains two parts: sensor-masked image generation and multi-sensor vision-language-action model architecture. The first part first extracts raw sensor data, then preprocesses it into 2D sensor images/heatmaps. After that, we use a semantic segmentation model to generate the masks from the RGB images following a VLM-generated prompt. Then we overlay the sensor images on the masked regions of RGB images to output sensor-masked images, which are the input for our multi-sensor vision-language-action model. The second part is a multi-sensor vision-language-action model backbone that is designed for sensor-masked images to capture each sensor’s input and avoid requiring a large amount of sensor data for training. We utilize the existing frozen vision encoders to encode sensor-masked images. For each sensor modality, we use individual multi-layer perceptron layers to align sensor image tokens with language and RGB image tokens. The tokens are concatenated together with language tokens as input for the large language model in the architecture, and then we generate the final action predictions using the action expert.
Figure 3: Sensor Data Processing Illustration. We propose a general sensor data processing pipeline applicable to various sensors, including (a) thermal camera, (b) mmWave radar, and (c) acoustic microphone array, by overlaying sensor information on top of RGB images as VLA model input. We update the prompt input to the SAM2 model when the task begins and then asynchronously check for updates, so that VLM delay does not affect the real-time processing of sensor-masked images.
@@ III-B Sensor Data Representation
The first part of our work is sensor-masked image generation, which includes sensor data preprocessing, segmentation mask processing, and sensor-masked image blending.
Preprocessing. We convert all raw measurements into a camera-like 2D spatial representation. Thermal camera already outputs raster images—infrared intensity, defined over image coordinates $(u,v)$. In contrast, the mmWave radar and the microphone array provide complex signal samples per array element, $x_{i,k}=A_{i,k}e^{j\psi_{i,k}}$ for the $k$-th element of array $i$, where $i\in\{\text{mmWave},\text{acoustic}\}$. To obtain inputs with consistent spatial mapping (horizontal and vertical viewing axes on a 2D grid), we compute azimuth–elevation heatmaps via conventional delay-and-sum beamforming [39]:
@@
where $\theta$ is the azimuth angle, $\phi$ is the elevation angle, $\lambda_{i}$ is the wavelength received by sensor $i$, $(x_{i,k},y_{i,k})$ is the position for $k^{th}$ array element (antenna or microphone) in array $i$. One example is shown in Fig. 1. Similar to the principle of RGB camera, the azimuth-elevation heatmap reveals the environmental information in a direct way for human understanding and acts as the sensor images in the following steps.
Segmentation and overlay. The next step is generating segmentation masks based on the task description and RGB images. We first send the task request text and RGB image to a vision-language model, GPT-4o [16], to generate the segmentation prompts that describe objects related to the task in the current scene, such as ‘red block/drink’, ‘black phone’, and ‘cardboard boxes’. These prompts are initialized when the task begins and are updated periodically to accommodate scene dynamics, such as the introduction of new objects. We run these updates in the background at a lower frequency than the robot’s primary action-generation loop, so that VLM output delay does not affect real-time robot manipulation actions. Then we input the segmentation prompts and RGB image into a segmentation model to generate image masks for task-related objects. Specifically, we use Grounded SAM 2 [33], which combines segment anything model [32] and open-set object detection model, grounding DINO [25]. The final output is a 0-1 matrix, where ‘1’ labels the masked regions for objects. The second step can be expressed as:
@@
$\text{mask}=\mathrm{SAM2}(l,\mathrm{I}_{\mathrm{RGB}}),$
where $l$ represents the segmentation prompts generated by VLM, $\mathrm{T}_{\text{task}}$ is the input task description, and $\mathrm{I}_{\mathrm{RGB}}$ is the RGB image. To generate the final sensor-masked images, we first spatially align each sensor modality with the RGB frame through a one-time calibration involving rotation and cropping. While high-precision pixel matching is not strictly required for the model to learn effectively, this ensures the sensor heatmaps are roughly centered on their physical counterparts. We then isolate the regions of interest by applying the masks generated in the previous step. This fusion process is formalized as:
@@
where $i\in\{\text{mmWave},\text{acoustic},\text{thermal}\}$ and $\alpha{\color[rgb]{0,0,0}\in[0,1]}$ is the hyper-parameter for image blending. The higher $\alpha$ is, the more sensor information remains for comprehensive understanding, but less correlation between the masked part and the unmasked RGB image remains. We set $\alpha$ as 1 by default based on our empirical testing. These consolidated sensor images then serve as the unified input for our multi-sensor VLA backbone.
@@ III-C Model architecture and training
The second part is the architecture design and training strategy of our multi-sensory vision-language-action model.
Existing VLA models typically encode RGB inputs with a vision encoder, followed by a multi-layer perceptron (MLP) to generate image tokens. Simultaneously, the task description is processed by a language tokenizer. These tokens are concatenated and fed into a large language model (LLM) backbone, which generates action tokens for robot manipulation. We reuse and adapt this architecture design.
We present a generalizable and efficient multi-sensory VLA model architecture design as shown in Fig. 2. We enable effective sensor data understanding by feeding sensor-masked images to the existing vision encoder. Then for each sensor modality, we input the encoded results into individual multi-layer perceptron (MLP) modules for projection to align with the language tokens. The projected tokens for each sensor are concatenated together with language tokens from the task description after passing through a language tokenizer. The LLM backbone takes these concatenated tokens as input. To generate the final robotic actions, we follow the architecture used in SmolVLA [34] and $\pi_{0}$ [3], where the LLM backbone is combined with a diffusion or flow matching based Action Expert for robotic action generation. We utilize conditional flow matching or diffusion within the action expert to model the continuous distribution of actions, which is inherited from the chosen base model. This allows the model to generate a full action chunk step by step from random noise. We use SmolVLA as our base model by default, modify its architecture, and train our model from its pretrained weights. Our approach is potentially compatible with various VLA backbones because our sensor-masked image representation allows processing sensor data as standard visual tokens.
The overall data flow can be described with:
@@
where $m$ is the number of sensors we use, $\mathrm{E}_{I}$ is the image encoder, and $\mathrm{E}_{L}$ is the language embedding layer, $\text{t}_{i}$ is the embedding of $i$th sensor image. We underline that our architecture does not require all sensors shown here. Instead, we allow flexible sensor setup according to the deployment scenario, budget, etc, for example using a single sensor like an infrared camera. This design is a general and flexible framework to process beyond-RGB sensor-masked images, compatible with existing Vision-Language-Action models. It means we can utilize the vision-language capability in existing model’s pretraining and understand unique features from each sensor easily with a few demonstration data.
During the model training, we start from the pretrained base model weight. We freeze the vision encoder, typical in various Vision-Language-Action models [34, 3, 17] and Vision-Language models finetuning [24, 10]. Meanwhile, we set all other weights trainable for multi-sensor perception in robotic manipulation. We initialize the individual MLP projectors for each sensor using weights from the pre-trained RGB projection layer of the base VLA model. This strategy leverages established visual feature mappings as a strong prior, allowing the model to adapt quickly to the unique sensor-masked images. Finally, we co-fine-tune these sensor-specific MLP modules alongside the unfrozen backbone weights using the collected demonstration datasets.